OpenStudy (anonymous):

Solve equation and write in simplest form. 7x/8=21 A.24 B. 18.375 C. 1.857 D. 19

OpenStudy (anonymous):

7x/8 = 21. Multiply both sides by 8 7x = 168 Divide both sides by 7 x = 24

OpenStudy (anonymous):

7x/8 = 21 7x = 21*8 7x = 168 x = 24